12/(x-4) - 5/(x+4) = 1/(x+4)
Answer by ilana(307)

First, we can look at the denominators and combine terms with equal denominators.
So we can say 12/(x-4) = 1/(x+4) + 5/(x+4).
Now, we can add those terms with the same denominator:
12/(x-4) = 6/(x+4).
Now, we have equivalent fractions, so we can cross multiply to get the equation:
12(x+4) = 6(x-4). We can distribute 6 and 12 to get 12x+48 = 6x-24.
Solving for x gives 6x=24, so x=4.
But we know x-4 in the denominator means x can't be 4, so this problem has no solution.